# Filtering with Confidence: When Data Augmentation Meets Conformal Prediction

### v1

@@ 4.2 Imbalanced Classification: tabular data examples
Table 1: Results of imbalanced classification: predictive performance metrics and data diversity (Stable Rank) averaged over 10 different splits. Dataset sizes (N), feature dimensions (d), and imbalance rates (Imb.) are given in parentheses. Higher values are better for all metrics presented. The best value is bolded and the second best value is underlined. Across all benchmarks, our CondCP-Filter consistently attains the best $F_{1}$ and increase data diversity, as reflected by higher stable rank.
In imbalanced classification, models often default to predicting the majority class, yielding misleadingly high accuracy while missing rare but critical events. For example, in the European Credit-Card Fraud dataset [^1] https://www.kaggle.com/datasets/mlg-ulb/creditcardfraud/data (0.17% frauds), labeling all cases as “non-fraud” achieves 99.8% accuracy but detects no fraud (He & Garcia, 2009; Japkowicz & Stephen, 2002). Data augmentation seems therefore a promising way of enhancing recall whilst maintaining precision.
We evaluate our method on three benchmark datasets spanning different imbalance regimes: European Credit-Card Fraud (Kaggle), Thyroid (OpenML), and MNIST-7 vs Others (OpenML). See the details of the dataset and experiment setup in Appendix D.7. In these settings, to generate new data, we train a Variational AutoEncoder (VAE) (Kingma & Welling, 2013; Sohn et al., 2015) to increase the number of samples from the minority class. Since gold-standard quality measures are not available in this setting, we use the procedure detailed in section 3.3, and use for our surrogate scores $\hat{A}$ a gradient boosting predictor, trained to predict the surrogate measure $\tilde{A}$. For the experiments presented in this subsection, $\tilde{A}$ is defined as the geometric mean of a $k$-nearest-neighbor similarity (to measure closeness to real minority data) and a cosine similarity (directional closeness to the reference data).
Table 1 reports $F_{1}$, precision, recall, and Stable Rank across five benchmarks. On severely imbalanced datasets such as credit-card fraud, quality-controlled augmentation clearly dominates both the unaugmented baseline and SMOTE; while SMOTE boosts recall, it inflates false positives, lowering precision. Our $\widehat{A}-$Filter and CondCP-Filter maintain recall while improving precision, yielding the best $F_{1}$. On moderate imbalance (Thyroid), all methods perform similarly, but our filters still outperform baselines and increase Stable Rank, indicating genuine diversity rather than duplication. For MNIST-7, where the signal is strong, unfiltered augmentation already works well; nonetheless, CondCP-Filter achieves the highest $F_{1}$ and superior precision–recall balance, showing the benefit of targeted acceptance even in easier tasks.
Beyond predictive performance, we also study diversity metrics of the training sets after augmentation with filtering. In particular, we compute the stable rank of the feature matrix, which captures the effective dimensionality of the sample cloud (Tsitsulin et al., 2023) . Whereas simple augmentation often inflates data density along a few dominant directions (due to interpolation), our method introduces genuinely new modes in the minority manifold, reflected in higher stable rank. These results indicate that quality-controlled generation is not only effective for balancing datasets, but also enhances geometric richness in ways that may improve generalization.
In this example, we wish to evaluate the performance of our method in a low-data regime for classification, with inputs of mixed qualities (with a distribution of 50% good inputs, 50% bad). In this setting, we expect the filtering procedure to be particularly useful.
@@ 4.3 Low-Data Regime with mixed-quality generations: an image analysis example
$\lambda$ Regime Training Test
Set Size Accuracy
- Unaugmented 172.8 $\pm$ 5.8 0.786 $\pm$ 0.023
- Unfiltered 3630.8 $\pm$ 317.4 0.802 $\pm$ 0.006
0.80 $\hat{A}$ Filter 897.0 $\pm$ 986.9 0.804 $\pm$ 0.036
0.75 CondCP Filter 1916.6 $\pm$ 158.4 0.823 $\pm$ 0.018
0.50 MargCP Filter 1334.8 $\pm$ 463.2 0.763 $\pm$ 0.055
Figure 4: Performance of the methods on the ImageNet dataset, averaged over 5 iterations.
5 Conclusion
Ethics statement
All authors confirm that this work adheres to the ICLR Code of Ethics. This research does not involve human subjects, personally identifiable information, or sensitive data. The datasets used are publicly available and have been properly cited. No potentially harmful applications, discriminatory outcomes, or security/privacy risks are anticipated from this study.
We have made efforts to ensure fairness, transparency, and integrity in both the methodology and the interpretation of results. All analyses and conclusions comply with ethical standards for reproducibility, research documentation, and integrity. No conflicts of interest or sponsorship influenced this work.
Appendix A The Use of Large Language Models (LLMs)
Appendix B Related Literatures
Data Augmentation
In Ding et al. (2024), LLM-based data augmentations are categorized into four categories: data creation, data reformation, data labeling and Human-LLM co-annotation. In this work we focus on the data reformation, which transforms existing data to produce new data. People have proposed data reformation approaches prior to the advent of pre-trained generative models, with the majorityrity of them being rule-based methods (Feng et al. (2021)). For instance, Easy Data Augmentation (EDA) Wei & Zou (2019) applies token-level perturbations like synonym replacement, random insertion, deletion, and swapping; Machine back-translation involves translating the original sentences into another language and then translating them back to the original language Sennrich et al. (2015); Edunov et al. (2018). Model-based methods, by contrast, leverage generative models to synthesize new text. Common examples include paraphrasing Kumar et al. (2019), semantic text exchange Feng et al. (2019) and masked word prediction followed by replacement Ng et al. (2020). The goal is to generate synthetic data that introduces diversity while maintaining semantic consistency (often referred to as label-preserving in classification problems (Xie et al., 2020)). Ideally, augmented data should not be too similar to the original (which limits diversity) nor too dissimilar (which risks domain shift and degraded performance).
Despite the ease of implementation, LLM-generated synthetic data is often noisy and distributionally different from the original data, potentially hindering model training (Zhang et al., 2022). To avoid this, some approaches focus on prompt engineering to guide generation Veselovsky et al. (2023); Gupta et al. (2023), some model-based data augmentation approaches directly estimate a generative process from the training set and sample from it (Anaby-Tavor et al., 2020), and some retrieval-based techniques improve the expressive performance of LLMs in data augmentation by introducing external knowledge (Chai et al., 2025). These approaches improve the quality of the generations, but still, there could be low-quality generations that ideally we would like to filter out. The filtering-based methods evaluate typically the generations based on quality metrics, such as human evaluation (Wang et al., 2021; Liu et al., 2022)(which can be expensive), similarity to the original input in paraphrasing Li et al. (2024b), confidence of LLM , or classifiers trained to distinguish real from synthetic data Veselovsky et al. (2023). Most of these methods either explicitly or implicitly leverage the prediction on the quality of the generations, which could be problematic when the prediction is not accurate.
Filtering.
Filtering methods are commonly based on the following strategies (Chai et al., 2025):
In practice, many augmentation pipelines combine multiple filters; for example, a heuristic may first remove obviously poor outputs, and then the top-$k$ most similar examples to the ground truth are retained (Chai et al., 2025). The goal is to balance fidelity and diversity: overly strict filters yield safe but low-diversity augmentations, while overly permissive filters risk introducing label noise or factual errors. Recent methods explicitly address this trade-off. For example, Mask-then-Fill (Gao et al., 2023) reports that infilling achieves a balance between novelty and distributional similarity to the source, likely through careful tuning of mask size and model parameters. In contrast, M4DA (Yao et al., 2024) promotes diversity by masking tokens to increase variance and then selecting variants with the highest semantic complexity. While the generated text must still preserve the original meaning, this preference for more complex rephrasings can yield stronger augmentation effects. Experiments on text classification benchmarks show that such methods can outperform conservative approaches, suggesting that filtering should not always default to the safest outputs—some controlled complexity, when consistent, is beneficial.
Appendix C Related Works On Conformal Prediction
Conformal Prediction
Given a dataset $\{(X_{i},Y_{i})\}_{i=1}^{N}$, a pretrained-predictor $h$ and a new text input $X_{n+1}$, conformal prediction (Vovk et al., 2005) attempts to construct a prediction $\hat{C}(X_{n+1})$ such that $\mathbb{P}\left(Y_{n+1}\notin\hat{C}(X_{n+1})\right)\leq 1-\alpha$ for some user-specified $\alpha$. Conformal prediction has the distribution-free property and it is finite-sample valid under the exchangeability of the data points $\{(X_{i},Y_{i})\}_{i=1}^{n+1}$. For instance, in split conformal prediction, one can define $S_{i}=\|Y_{i}-h(X_{i})\|$ and then set $\hat{C}(X_{n+1})=\{y:\|y-h(X_{n+1})\|\leq\tau\}$ where $\tau=\text{quantile}(\{S_{i}\}_{i=1}^{n}\cup\{\infty\},1-\alpha)$ . This type of method provides a guarantee on marginal coverage. Previous studies have demonstrated that achieving exact conditional coverage is impossible without any further distributionally assumption (Barber et al., 2021). Nevertheless, researchers have developed methods to achieve conditional coverage with controllable error rates (Zhang & Candès, 2024; Gibbs et al., 2025).
Conformal Prediction and LLMs
Researchers have increasingly explored the application of conformal prediction (CP) frameworks in generative models for factuality control, motivated by CP’s ability to provide distribution-free inference. In (Ren et al., 2023; Kumar et al., 2023), CP is employed to identify probability thresholds for next-token generation, thereby selecting response candidates. Several works have proposed CP methods that do not require access to model logits (Su et al., 2024). For instance, (Shahrokhi et al., 2025; Quach et al., 2023) use CP to determine the number of generations needed to construct a prediction set that includes at least one truthful response or satisfies a specified confidence level. Other approaches, such as (Mohri & Hashimoto, 2024; Cherian et al., 2024), segment LLM outputs into individual claims and apply CP to select factual ones. Additionally, Gui et al. (2024) extends CP to multiple test units with a focus on ensuring valid false discovery rate (FDR) control. Despite the successes in these applications, how CP can be applied in data augmentation is under-explored, perhaps due to its unsupervised nature.
Conditional Conformal Prediction
$\mathbb{P}\big(Y_{n+1}\in\hat{C}(X_{n+1})\mid X_{n+1}=x\big)=1-\alpha,$
which may differ significantly across inputs. This limitation is critical in sensitive applications (e.g., medicine, finance), where systematic under-coverage on certain subgroups undermines reliability. Prior work shows that in distribution-free settings, exact conditional coverage is impossible: any set satisfying it must degenerate to $\hat{C}(X_{n+1})=\mathbb{R}$ with infinite expected size (Barber et al., 2021).
C.1 Additional details on conditional conformal
In our setting, the conformity score $S_{i_{0}}$ is unknown for every $i_{0}\in\mathcal{D}_{\mathrm{aug}}$; we therefore impute a value $S$ for each such test index and solve a single regularized quantile problem that treats the imputed test pair symmetrically with the calibration data. Following Gibbs et al. (2025), we estimate a high–probability upper bound for these scores $\{S_{i}\}_{i\in\mathcal{D}_{\mathrm{calib}}}\cup S$ by fitting a regularized kernel quantile regression:
$\hat{f}_{S}\;=\;\arg\min_{f\in\mathcal{F}^{*}}\Bigg\{\frac{1}{|\mathcal{D}_{\mathrm{calib}}|+1}\sum_{i\in\mathcal{D}_{\mathrm{calib}}}\!\ell_{\alpha}\!\big(S_{i}-f(X_{i})\big)\;+\;\frac{1}{|\mathcal{D}_{\mathrm{calib}}|+1}\,\ell_{\alpha}\!\big(S-f(X_{i_{0}})\big)\;+\;\frac{\gamma}{2}\,\|f_{W}\|_{W}^{2}\Bigg\},$ (5)
By the representer theorem (Kimeldorf & Wahba, 1971), the optimizer admits the finite expansion
with coefficient vector $\hat{\upsilon}_{S}\in\mathbb{R}^{|\mathcal{D}_{\mathrm{calib}}|+1}$ and intercept $\hat{\beta}_{S}\in\mathbb{R}$. Accordingly, the fitted RKHS component is of form $\hat{f}_{W}(x)=\frac{1}{\gamma}\sum_{i\in\mathcal{D}_{\mathrm{calib}}\cup\{i_{0}\}}\hat{\upsilon}_{S,i}\,W(x,X_{i})$ As shown in the discussion of Anonymous (2026), the coefficients $\hat{\upsilon}_{S}$ depend affinely on the imputed value $S$ and the mapping $S\mapsto\hat{\upsilon}_{S}$ is nondecreasing. Consequently, the event $S\leq\hat{f}_{S}(X_{i_{0}})$ is equivalent to the linear inequality $\hat{\upsilon}_{S,i_{0}}\leq 1-\alpha$. Following the standard randomized conformalization in Anonymous (2026); Gibbs et al. (2025), we replace $1-\alpha$ by a draw $U\sim\mathrm{Unif}(-\alpha,\,1-\alpha)$, and define the final fitted cutoff by
$\hat{s}_{i_{0}}\;=\;\max\{\,S:\ \hat{\upsilon}_{S,i_{0}}\leq U\,\}.$ (7)
Coverage guarantee.
Lemma 2 (Coverage; cf. Gibbs et al. (2025); Cherian et al. (2024))
$\mathbb{E}\!\left[f(X_{i_{0}})\left\{\mathbf{1}\!\left(\mathcal{L}_{\lambda}\big(\mathcal{S}(\widehat{\mathbf{A}}_{i_{0}};\hat{s}_{i_{0}}),\,\mathbf{A}_{i_{0}}\big)\leq\rho\right)-(1-\alpha)\right\}\right]\;=\;-\gamma\,\mathbb{E}\!\left[\,\big\langle\hat{f}^{\;\hat{s}_{i_{0}}}_{W},\,f_{W}\big\rangle_{W}\,\right].$
C.2 Proof of Lemma 1
For the localized conformal prediction, we adapt Lemma 2 to a class of covariate shifts induced by the density kernel $W$. Under the setting of Lemma 1, the tuples
$(X_{1},\{G_{1k}\}_{k\in[K]},Y_{1}),\dots,(X_{n},\{G_{nk}\}_{k\in[K]},Y_{n})$
$X_{i_{0}}\sim P_{X},\qquad Y_{i_{0}}\mid X_{i_{0}}\sim P_{Y\mid X},\qquad(G_{i_{0}k})_{k=1}^{K}\mid(X_{i_{0}},Y_{i_{0}})\sim h(X_{i_{0}},Y_{i_{0}},\tau),$
For any realization $x^{\prime}\in\Omega$, Bayes’ rule yields
$\big(X_{i_{0}},(G_{i_{0}k})_{k=1}^{K},Y_{i_{0}}\big)\,\big|\,X^{\prime}_{i_{0}}=x^{\prime}\penalty 10000\ \sim\penalty 10000\ \frac{W(x,x^{\prime})}{\mathbb{E}[\,W(X,x^{\prime})\,]}\;dP_{(X,\mathbf{G},Y)}(x,\mathbf{G},y),$
$\mathbb{E}\!\left[\mathbf{1}\!\left\{\mathcal{L}_{\lambda}\!\big(\mathcal{S}(\widehat{\mathbf{A}}_{i_{0}};s_{i_{0}}),\mathbf{A}_{i_{0}}\big)\leq\rho\right\}-(1-\alpha)\,\middle|\,X^{\prime}_{i_{0}}=x^{\prime}\right]$
$\qquad=\;\frac{\mathbb{E}\!\left[W(X,x^{\prime})\cdot\Big(\mathbf{1}\!\left\{\mathcal{L}_{\lambda}\!\big(\mathcal{S}(\widehat{\mathbf{A}}_{i_{0}};s_{i_{0}}),\mathbf{A}_{i_{0}}\big)\leq\rho\right\}-(1-\alpha)\Big)\right]}{\mathbb{E}[\,W(X,x^{\prime})\,]}$
@@ C.2 Proof of Lemma 1
$\qquad=\;\frac{-\,\gamma\,\mathbb{E}\!\left[\,\hat{f}^{\;\hat{s}_{i_{0}}}_{W}(x^{\prime})\,\right]}{\mathbb{E}[\,W(X,x^{\prime})\,]}.\quad$
Using the finite expansion of the fitted RKHS component,
we can further write
$\mathbb{E}\!\left[\mathbf{1}\!\left\{\mathcal{L}_{\lambda}\!\big(\mathcal{S}(\widehat{\mathbf{A}}_{i_{0}};s_{i_{0}}),\mathbf{A}_{i_{0}}\big)\leq\rho\right\}-(1-\alpha)\,\middle|\,X^{\prime}_{i_{0}}=x^{\prime}\right]$ $=\;\frac{-\,\mathbb{E}\!\left[\sum_{i\in\mathcal{D}_{\mathrm{calib}}\cup\{i_{0}\}}\hat{\upsilon}_{\hat{s}_{i_{0}},i}\,W(X_{i},x^{\prime})\right]}{\mathbb{E}[\,W(X,x^{\prime})\,]},$
which completes the localized reweighting conformal prediction.
Appendix D Experimental details
D.1 Clinical Text Classification
Prompt for Data Generation
You are given a description of a disease.
Description: {symptom}
Task: Extend the symptom description with additional details that
still plausibly describe the SAME disease.
- Write EXACTLY 5 sentences.
- Do not copy wording from the original; paraphrase and add
plausible details consistent with the same condition.
- Avoid lists, bullets, headings, or numbering; just 5 full
sentences in a single paragraph.
- No disclaimers, no citations, no markdown.
Prompt for Evaluation
You are evaluating individual symptom descriptions for diseases.
Scoring instructions:
- Assign each description a score between 0 and 1, rounded to two
decimal places.
- Criteria: The description should plausibly match the specified
disease and avoid confusion with other diseases.
- Use the full 01 range: 1 = perfectly clear, specific, and
accurate; 0 = completely unusable.
- 0.5 is the threshold: any description with a score <= 0.5 should
be dropped to prevent misclassification.
For reference, here is the complete list of possible diseases:
{disease_ls}
Output requirements:
- Output only the scores, one per line, in the same order as the
input  cases.
- Do not include explanations, text, or formatting other than the
numeric scores.
Case : Disease: {diag} : Symptom: {symp}
D.2 Abstract Topic Classification
$S(\hat{\mathbf{A}}_{i},\mathbf{A}_{i})=\inf\left\{\tau:\big|\{G_{ik}:1\leq k\leq 3,\,\hat{A}_{ik}\geq\tau,\,A_{ik}<0.5\}\big|=0\right\}.$
Prompt for Data Generation
You are given a statistical abstract.
Abstract: {abstract}
Task: Extend the abstract with additional details that remain
consistent with the SAME statistical topic.
- Write EXACTLY 6 sentences.
- Do not copy wording from the original; paraphrase and add
plausible extensions consistent with the same subject.
- Avoid lists, bullets, headings, or numbering; just 6 full
sentences in a single paragraph.
- No disclaimers, no citations, no markdown.
Prompt for Evaluation
You are evaluating individual sentences from extended statistical
abstracts.
Scoring instructions:
- Assign each sentence a score between 0 and 1, rounded to two
decimal places.
- Criteria: The sentence should plausibly match the specified topic,
remain coherent, and avoid drifting into other topics from the list.
- Use the full 01 range: 1 = perfectly clear, on-topic, and
informative; 0 = completely unusable.
- 0.5 is the threshold: any sentence with a score <= 0.5 should be
dropped to prevent topic drift.
Output requirements:
- Output only the scores, one per line, in the same order as the
input cases.
- Do not include explanations, text, or formatting other than the
numeric scores.
Case : Topic: {topic} : Sentences: {sent}
D.3 Twitter Message Sentiment Analysis
$S(\hat{\mathbf{A}}_{i},\mathbf{A}_{i})=\inf\left\{\tau:\big|\{G_{ik}:1\leq k\leq 5,\,\hat{A}_{ik}\geq\tau,\,A_{ik}<0.5\}\big|\leq 1\right\}.$
Prompt for Data Generation
You are given a short Twitter message.
Message: {tweet}
Task: Extend the message with additional content that preserves the
SAME sentiment and topic.
- Write EXACTLY 5 sentences.
- Paraphrase and expand naturally; do not copy wording from the
original.
- Vary phrasing, tone, and detail while remaining consistent with
the sentiment.
- Avoid lists, bullets, hashtags, mentions, links, or numbering;
produce 5 full sentences in a single paragraph.
- No disclaimers, citations, or markdown.
Prompt for Evaluation
You are evaluating individual sentences for sentiment consistency.
Scoring instructions:
- Assign each sentence a score between 0 and 1, rounded to two
decimal places.
- Criteria: The sentence should clearly reflect the SPECIFIED
sentiment, remain coherent, and avoid conflicting emotions.
- Use the full 01 range: 1 = perfectly consistent and natural;
0 = completely unusable.
- 0.5 is the threshold: any sentence with a score <= 0.5 should
be excluded.
Output requirements:
- Output only the scores, one per line, in the same order as the
input cases.
- Do not include explanations, text, or formatting beyond the
numeric scores.
Case: Sentiment: {senti} Sentence: {sent}
D.4 Diversity of Augmented Text
To evaluate the diversity of the augmentations, we compute the Shannon entropy of different augmentations (see Table 2). We find that augmentation generally increases the diversity of the training data. The filtered versions typically show lower diversity compared with the unfiltered versions, since poor-quality generations are excluded. In particular, the CondCP-filtered generations exhibit lower diversity than those filtered by other algorithms. This outcome is intuitive given the nature of conformal prediction. Nonetheless, the reduction in diversity is relatively small, highlighting that the CondCP filter effectively preserves the essential information contained in the training data.
Diagnosis 6.92 8.52 8.14 8.14 8.02
Abstract 9.64 9.83 9.81 9.81 9.76
Sentiment Analysis 8.69 9.34 9.23 9.23 8.96
D.5 Comparision between Gemini-2.5-Pro scores and Gemini-2.5-Flash scores
D.6 Examples of metrics
Data Type Similarity Metric Description References
Text BERT-based Cosine Similarity Semantic coherence via embeddings Devlin et al. (2018); Zhang et al. (2020)
METEOR Harmonic mean of precision and recall Banerjee & Lavie (2005)
Sentence-BERT Sentence-level semantic similarity Reimers & Gurevych (2019)
Image CLIP Score Semantic alignment via text-image embeddings Radford et al. (2021)
Structural Similarity Index (SSIM) Structural similarity of images Wang et al. (2004)
Fréchet Inception Distance (FID) Distance in feature distributions Heusel et al. (2017)
Peak Signal-to-Noise Ratio (PSNR) Pixel-level error measurement Hore & Ziou (2010)
Learned Perceptual Image Patch Similarity (LPIPS) Learned perceptual similarity Zhang et al. (2018)
Table 3: Examples of affinity metrics used in our filtering framework.
D.7 Experiment Details for Imbalanced Classification
@@
We generate minority samples with a Conditional Variational Autoencoder (CVAE) (Kingma & Welling, 2013; Sohn et al., 2015) that is conditioned on an actual minority seed. Let $h_{\psi}:\mathbb{R}^{d}\to\mathbb{R}^{c}$ be a small context net that maps a reference minority instance $x$ to a context vector $c=h_{\psi}(x)$. The encoder receives the concatenation $[x,c]$ and outputs a Gaussian $q_{\phi}(z\mid x,c)=\mathcal{N}(\mu_{\phi}(x,c),\operatorname{diag}(\sigma^{2}_{\phi}(x,c)))$; the decoder reconstructs $x$ from $[z,c]$ via $p_{\theta}(x\mid z,c)$. Both encoder and decoder are two-layer MLPs with ReLU activation.
@@
We apply conditional conformal filtering that operates in a learned latent representation of the data (Anonymous, 2026). Specifically, we project the feature space into a lower-dimensional latent embedding using Principal Component Analysis (PCA) before applying the conformal calibration step. For each dataset, we tune the latent dimension to reflect its scale: 2 for Thyroid and Credit Card Fraud, and 16 for MNIST-7.

### v2

@@ D.10 Additional Experiments on Imbalanced Classification
MCI (RAZRIV) (N=1,700, d=111, Imb.=3.3%) Unaugmented 0.033 $\pm$ 0.070 0.200 $\pm$ 0.422 0.018 $\pm$ 0.038 26.886 $\pm$ 0.635
@@
$\widehat{A}-$Filter 0.182 $\pm$ 0.048 0.111 $\pm$ 0.029 0.500 $\pm$0.150 12.142 $\pm$ 1.758
CondCP-Filter 0.211 $\pm$ 0.067 0.155 $\pm$ 0.048 0.400 $\pm$ 0.143 19.228 $\pm$ 5.291
D.10.1 Experiment Details
@@
  • Myocardial infarction complications (UCI-579): A clinical dataset comprising 1,700 patient records. We define a binary classification task targeting RAZRIV (myocardial rupture), with a positive class prevalence of approximately 3.3%. Features include 111 clinical attributes collected at admission and during the first three days of hospitalization, such as demographics, anamnesis, ECG characteristics, and laboratory results.
@@
We generate minority samples with a Conditional Variational Autoencoder (CVAE) (Kingma and Welling, 2013; Sohn et al., 2015) that is conditioned on an actual minority seed. Let $h_{\psi}:\mathbb{R}^{d}\to\mathbb{R}^{c}$ be a small context net that maps a reference minority instance $x$ to a context vector $c=h_{\psi}(x)$. The encoder receives the concatenation $[x,c]$ and outputs a Gaussian $q_{\phi}(z\mid x,c)=\mathcal{N}(\mu_{\phi}(x,c),\operatorname{diag}(\sigma^{2}_{\phi}(x,c)))$; the decoder reconstructs $x$ from $[z,c]$ via $p_{\theta}(x\mid z,c)$. Both encoder and decoder are two-layer MLPs with ReLU activation.
@@
We apply conditional conformal filtering that operates in a learned latent representation of the data (Jung et al., 2025). Specifically, we project the feature space into a lower-dimensional latent embedding using Principal Component Analysis (PCA) before applying the conformal calibration step. For each dataset, we tune the latent dimension to reflect its scale: 2 for Thyroid and Credit Card Fraud, and 16 for MNIST-7.
D.10.2 Evaluating the quality of the (selected) generations
We examine the MNIST 7 example, an imbalanced classification task where MNIST digits are classified as 7 or not 7, with 7s being underrepresented. We evaluate the effect of temperature and selection on the diversity of the generated samples, as the results in this example should be easily interpretable visually.
Understanding the effect of temperature on the diversity of the samples.
Figure 8 highlights a few examples of generations of the digit $7$ for different temperatures $\tau$. In low temperature settings (e.g. $\tau=0.1$), the model generates almost identical samples. In moderately high temperature settings ($\tau=2$), the model starts to generate more variable shapes of the digit 7. However, as the temperature becomes too high ($\tau=10$), the synthetic data become extremely noisy.
Figure 9 and 10 further illustrate the temperature effect through principal component analysis, comparing real and generated data. As the temperature ($\tau$) increases, the synthetic samples gradually explore a wider area of the real data. However, excessively high temperatures (e.g., $\tau=10$) cause the generator to sample outside the MNIST distribution, resulting in points that do not align with the original data’s structure.
Figure 8: Generated minority digit (7) by the VAE model for varying temperature ($\tau$).
Figure 9: PCA visualization of real MNIST-7 digits and VAE-generated samples under different temperature values $\tau$. The gray circle points denote the real data, and the blue triangular points denote the generated data.
Figure 10: PCA visualization of real samples from thyroid dataset and VAE-generated samples under different temperature values $\tau$. The gray circle points denote the real data, and the blue triangular points denote the generated data.
To quantify these effects and directly examine the role of temperature, we measure the diversity of the (unfiltered) generated samples for different values of the parameter $\tau$ using the stable rank. The results, shown in Table 8, confirm that increasing the sampling temperature produces higher-diversity synthetic data; however, but this diversity is uncontrolled. Beyond moderate temperatures, the generator begins to sample outside the real data manifold, producing overly noisy or implausible examples. Figure 8 illustrates this phenomenon, where high-temperature samples turn out to be overly noisy. This supports the motivation stated in the introduction: while higher temperatures can, in principle, expose rare modes, naively relying on high-temperature augmentation is harmful because it injects low-quality, out-of-support samples.
Table 8: Diversity measure by Stable Rank of unfiltered augmentation with varying temperature ($\tau$). Mean and standard deviation computed across different splits (seed). The smaller $\tau$ will lead to generate the samples closer to the existing point, and the higher $\tau$ will lead to more noisy generation.
$\tau$ Thyroid Credit Card Fraud MNIST 7
0.1 6.072 $\pm$ 0.689 1.960 $\pm$ 0.039 12.048 $\pm$ 0.754
1 8.238 $\pm$ 0.773 1.989 $\pm$ 0.034 12.020 $\pm$ 0.745
2 8.026 $\pm$ 1.075 1.998 $\pm$ 0.095 12.125 $\pm$ 0.707
Understanding the effect of selection.
Figures 11 and 12 show examples of accepted and rejected samples, respectively. As expected, we see that the rejected samples feature more low quality (extremely blurry and jagged) sevens, compared to the selected ones: the condCP selection seems to select samples that are more realistic.
Figure 11: Examples of accepted generations by CondCP with $\tau=0.1,\lambda=0.5,\rho=2$.
Figure 12: Examples of generations filtered out by CondCP with $\tau=0.1,\lambda=0.5,\rho=2$.
This effect extends to other datasets. In Figure 13, we visualize the effect of the selection on the Thyroid dataset. To this end, we first extracted important predictors using regression on training dataset, and the three variables, on thyroxine (binary treatment indicator), T3 (serum triiodothyronine level), and TT4 (total thyroxine level), emerged as significant predictors. The plots show, for each variable and each filtering strength $\lambda$, how the distribution of accepted synthetic samples aligns with the real minority-class distribution. Across features, accepted samples (blue) consistently match the true minority distribution (black) better than rejected samples (gray), demonstrating that the filtering criterion preferentially retains synthetic points that lie on the true data-support for the minority class.
Figure 13: Using the thyroid dataset, we first fit a simple logistic regression on 60% of the real data to predict the minority thyroid-disease class. Three variables, on thyroxine (binary treatment indicator), T3 (serum triiodothyronine level), and TT4 (total thyroxine level), emerged as significant predictors. The tolerance parameter $\rho$ is fixed to be 2.